# MINT: Multi-Vector Search Index Tuning

### v1

@@ 16                         $EK(i,cover)\leftarrow EK(i-1,cover-cvr)\cup\{ek_{i}:\max_{j\in cvr}ek_{i}\_j\}$
{$cost_{cover}(cover,x_{1})&\quad\text{when}~{}i=1,\\ \min_{cvr\subseteq cover}\Big{(}DP(i-1,cover-cvr)\Big{.}&\\ \Big{.}+cost_{cover}(cvr,x_{i})\Big{)}&\quad\text{when}~{}2\leq i\leq|X|.\end{aligned}\right.$
where $cost_{cover}(cvr,x_{i})=\max_{j\in cvr}cost_{idx}(q,x_{i},ek_{i}\_j)$ is the cost to cover item subset $cvr$ using index $x_{i}$. The space complexity is $O(2^{k^{\prime}}\cdot|X|)$ for keeping only $DP(i-1,cover),DP(i,cover),EK(i-1,cover)$, and $EK(i,cover)$. The time complexity of DP is $O(|X|\cdot 2^{2k^{\prime}})$, bringing the overall time complexity down to $O(k^{\prime}\cdot\sum_{x_{i}\in X}x_{i}.dim+|X|k^{\prime}\log k^{\prime}+|X|\cdot 2^{2k^{\prime}})$. In practice, we set $k^{\prime}$ to a small constant, say $5$, to get acceptable accuracy, and the overall time complexity becomes $O(\sum_{x_{i}\in X}x_{i}.dim+|X|)$, which is linear in the sum of dimensions and number of indexes. Algorithm 2 is the pseudo code for a sample $gt_{sp}(q)$.
In Mint, we use both Algorithm 1 and 2 in different conditions. When the number of indexes $|X|$ is no more than $3$, we apply Algorithm 1. Otherwise, we apply Algorithm 2 for better efficiency.
3.3.2. Cost Estimator and Recall Estimator
Here we explain how to estimate cost and recall for the proposed algorithms above.
Figure 5. Observed $numDist(q,x_{i},ek_{i})$ using (a) HNSW and (b) DiskANN indexes on four example datasets can be represented by linear models.
Figure 6. Observed recall using (a) HNSW and (b) DiskANN indexes on four example datasets can be represented by logarithmic models.
@@ 3.3.2. Cost Estimator and Recall Estimator
As we will see in Section 5, the cost and recall estimators perform reasonably well. The training data size is only 1% of the database and the training time is between 3 and 32 seconds. Note that our algorithms are flexible and can take any cost and recall estimators (Figure 1) if more accurate estimators become available.
4. Configuration Searcher
Configuration Searcher is the main module that searches for the best configuration to solve Problem 1. It explores various configurations, and for each promising configuration, it asks the Query Planner “if these indexes are created, what estimated cost and recall will the best query plan have” (similar to a what-if call in a relational database (Chaudhuri and Narasayya, 1998)). Next, we prove its NP-Hardness and propose a beam search algorithm.
4.1. NP-Hardness of Configuration Search
We prove the NP-Hardness of Configuration Search by reducing from the Densest $\mathscr{k}$-Subgraph problem (Chaudhuri et al., 2004; Feige and Seltser, 1997). We replace $\mathscr{k}$ using $\mathscr{g}$ hereafter to distinguish $\mathscr{k}$ from the top-$k$ for queries.
Problem 4 (Densest $\mathscr{g}$-Subgraph).
Given a graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$, find a subgraph $\mathscr{G}_{sub}=(\mathscr{V}_{sub},\mathscr{E}_{sub})$ with exactly $\mathscr{g}$ vertices such that the number of edges in the subgraph is maximal.
@@ 4.1. NP-Hardness of Configuration Search
The idea is to map each vertex to an index and each edge to a cost benefit of a query. When two vertices are in $\mathscr{G}_{sub}$, the query can benefit from the corresponding indexes.
@@
(Sketch) Let each vertex $\mathscr{v}^{i}\in\mathscr{V}$ be a single-column index $x^{i}$, each edge $(\mathscr{v}^{i},\mathscr{v}^{j})\in\mathscr{E}$ be a two-column query $q^{i,j}$ in Problem 1, each probability $p^{i,j}$ be the same constant, and each query find the same top-$k$. Then let $\mathscr{g}$ be $\theta_{storage}$, and $\theta_{recall}=100\%$. Finally, let
$latency(q^{i,j},X)=\left\{\begin{aligned} \small k&\quad\text{when}~{}\mathscr{v}^{i}\in\mathscr{G}_{sub}\textbf{~{}and~{}}\mathscr{v}^{j}\in\mathscr{G}_{sub},\\ \small\mathscr{L}&\quad\text{otherwise}.\end{aligned}\right.$
where $\mathscr{L}\gg k$ is a long latency of scanning almost the whole database. This latency function is possible in practice. Specifically, let each $q^{i,j}$ be a vector of all 0 but the $i$-th and $j$-th dimensions are 1. Let ground truth $gt(q^{i,j})=gt^{i}\cup gt^{j}$, where $|gt^{i}|=|gt^{j}|=k/2$. In $gt^{i}$, the vectors’ $i$-th dimension is 1 and other dimensions are 0. Similarly, $gt^{j}$ has vectors with only $j$-th dimension equal to 1. All other dummy vectors are $[0.1,...,0.1]$. So ground truth vectors have $score=1$ and dummy vectors have $score=0.2$. Plus, $gt^{i}$ rank at the top of $x^{i}$ but $gt^{j}$ rank near the bottom of $x^{i}$ (by $score$ on $\textbf{v}_{i}$), and $gt^{j}$ rank at the top of $x^{j}$ but $gt^{i}$ rank near the bottom of $x^{j}$ (by $score$ on $\textbf{v}_{j}$). As a result, when $\mathscr{v}^{i},\mathscr{v}^{j}\in\mathscr{G}_{sub}$, meaning $x^{i}$ and $x^{j}$ are both picked, the latency is $k$ by retrieving $gt^{i}$ from $x^{i}$ and $gt^{j}$ from $x^{j}$. Otherwise, we must scan other dummy items in latency $\mathscr{L}$.
Now each subgraph $\mathscr{G}_{sub}$ can be mapped to a configuration $X$ and vice versa. The total latency is $(|\mathscr{E}|-|\mathscr{E}_{sub}|)\cdot\mathscr{L}+|\mathscr{E}_{sub}|\cdot k=|\mathscr{E}|\mathscr{L}-|\mathscr{E}_{sub}|(\mathscr{L}-k)$. So maximizing $|\mathscr{E}_{sub}|$ is equivalent to minimizing the total latency and vice versa.
4.2. Beam Search Algorithm
Input: Query workload $W$; Constraints $\theta_{storage}$ and $\theta_{recall}$; Subset difference $di$; Seed limit $se$; Beam width $b$; Improvement threshold $im$; Query Planner.
@@ 4.2. Beam Search Algorithm
1
[…]
3
[…]
5
[…]
7
[…]
9
10repeat
11       $Config\leftarrow\varnothing$
12       foreach $X\in Best,x\in Candidate$ do
13             $X^{\prime}\leftarrow X\cup\{x\}$
14             Get plans for all queries using $X^{\prime}$ from Query Planner and remove unused indexes from $X^{\prime}$
15
16            if $X^{\prime}$ satisfies constraints $\theta_{storage}$ and $\theta_{recall}$ then
17                   $Config\leftarrow Config\cup X^{\prime}$
18
19
20      $Best\leftarrow$ $b$ lowest-cost configurations in $Config$
21
22until Cost improvement $\leq im$
23
24return lowest-cost configuration $X^{*}$ that we have ever seen
@@ 20      $Best\leftarrow$ $b$ lowest-cost configurations in $Config$
We use a beam search algorithm to find optimal configurations as shown in Algorithm 3. We begin by finding a set of candidate indexes from all queries (Line 3 and 3). For each query $q_{i}$, the candidate is a subset whose size is close to $q_{i}.vid$. For instance, when $|q_{i}.vid|=5$ and $di=2$, candidates of $q_{i}$ are subsets with $\geq 3$ columns. Then we create a set of seed configurations from all queries (Line 3). For each query $q_{i}$, the seed configurations are all subsets of $Cand(q_{i})$ with no more than $se$ indexes. Then we keep the best $b$ configurations in Line 3, and perform beam search from Line 3 to Line 3. Within each iteration, we try adding one more index to the existing configuration without violating constraints. The loop stops when the cost improvement is no more than parameter $im$.
@@
We implement several optimizations during search. For example, we may search more than $b$ seeds at the beginning to find better configurations if $|Candidate|$ is limited. We also cache the query plans of $(q,X)$ pairs. Instead of calling Query Planner for each $(q,X^{\prime})$ in Line 3, we first identify useful indexes $X^{\prime}_{q}=\{x~{}|~{}x\in X^{\prime}\text{~{}and~{}}x.vid\subseteq q.vid\text{~{}and~{}}|x.vid|\geq|q.vid|-di\}$. Then, if the query plan of $(q,X^{\prime}_{q})$ is cached, we use the cached plan without calling Query Planner. Similarly, to save running time, we cache and pass relevant $ek$ to Query Planner if $ek$ has been computed for a $(q,x)$ pair where $x$ is an index.
5. Experiments
In our experimental evaluation of Mint, we answer the following questions: Q1: What is the quality of indexes and plans generated by Mint for various workloads, compared with state-of-the-art solutions? (Section 5.2). Q2: What is the scalability of Mint? How is the runtime of Mint affected by key factors such as query number and storage constraint? (Section 5.3).
5.1. Setting
We implement the Configuration Searcher and Query Planner in Python, and the underlying indexes in C++. All experiments are conducted on a single machine equipped with an AMD EPYC 7763 64-core processor, 64GB of DDR4 memory, and running Ubuntu 24.04.1 LTS. We repeat each experiment with different random seeds five times and report the mean values.
Datasets
To evaluate the performance of Mint, we use a mix of semi-synthetic datasets and one real-world dataset (Table 1). Due to the limited availability of public large-scale multi-feature data and restrictions on the use of company datasets, we generate semi-synthetic datasets by combining single-feature embedding datasets including GloVe25 (Pennington et al., 2014), GloVe50, GloVe100, GloVe200, SIFT1M (Jégou et al., 2011), Yandex Text-to-Image (Simhadri et al., 2022), Deep1M (Yandex and Lempitsky, 2016), and Music (Morozov and Babenko, 2018). We sample an equal number of rows from each single-feature dataset and treat each as a separate column, mimicking the multi-feature scenario.
Our real-world dataset is News, consisting of 0.1M rows and four embeddings extracted from titles, descriptions, contents, and images of BBC News (Li et al., 2024). We get the embeddings for images and titles using the CLIP model (Radford et al., 2021), which handles images and short text with less than $77$ tokens. For longer text, such as news description and content, we use Nomic-BERT model (Nussbaum et al., 2024).
@@ 5.1. Setting
Table 1. Datasets, each of which is a column (i.e., feature).
Workload Type # rows # cols # queries Avg. # cols per query
Naïve Semi-Synthetic 1M 3 4 2
@@
News Real 0.1M 4 6 2.5
Table 2. Workloads.
Figure 7. Using HNSW, (A) Mint outperforms PerColumn on all four workloads, and Mint’s latency is close to the optimal PerQuery’s, although PerQuery uses 33% to 50% more storage. (B) The speedup ranges from 2.1×2.1$\times$2.1 × to 7.2×7.2$\times$7.2 ×. (C) Mint outperforms PerColumn because Mint has lower cost.
Workloads
Table 2 summarizes the workloads in our evaluation. The small semi-synthetic workload, Naïve, contains three columns (GloVe100, SIFT1M, and Yandex Text-to-Image) and 1 million rows, which is manually crafted for the case study. It has four queries. The larger semi-synthetic workloads, BiSimple and BiComplex, contain all eight semi-synthetic columns in Table 1 and 1 million rows. We generate each query following a binomial distribution: each column has a probability $\mathscr{p}$ of being included in a query. Intuitively, a smaller $\mathscr{p}$ generates simpler queries with fewer columns. We use $\mathscr{p}=0.3$ for BiSimple and $\mathscr{p}=0.5$ for BiComplex. Each workload has 12 queries. The News workload has six queries on the four real embeddings. The columns in each query also follow a binomial distribution with $\mathscr{p}=0.5$. In all workloads, the queries’ probabilities are drawn from a uniform distribution and normalized.
@@
We compare Mint against two baselines: PerColumn and PerQuery. PerColumn means one-index-per-column (Wang et al., 2021b; Zhang et al., 2023). It constructs an independent index for each column in the database. Given a workload, each query is answered by the indexes corresponding to the columns involved in the query. PerQuery means one-index-per-query. It creates one index for each query in the workload, so it violates the storage constraints in all workloads we experiment with. However, we still consider it to be a lower bound of latency. All approaches (Mint, PerColumn, and PerQuery) use two widely adopted ANN indexes, HNSW and DiskANN. In Mint, we sample 1% of the data to train the cost estimator and recall estimator (Section 3.3.2).
@@
We use cosine similarity as the distance measure in all experiments. We measure the latency and the cost as a proxy for latency (Section 3.1). Given a workload of queries with probabilities, we measure the probability-weighted average latency and cost across queries.
@@
The storage in bytes is proportional to the number of indexes when a given workload has a fixed number of items, and our maximum degree of HNSW and DiskANN defaults to $16$. So we use the number of indexes as the storage.
The default storage constraint is the number of columns because PerColumn requires one index per column. As Table 2 shows, the storage constraints of Naïve, BiSimple, BiComplex, and News are $3,8,8$, and $4$, respectively. We apply the default storage to PerColumn and Mint for a fair comparison. PerQuery requires one index per query and uses $4,12,12$, and $6$ storage, which violates the constraint but is shown as a reference. In addition, since Mint is designed to be flexible with storage, we evaluate Mint under various storage constraints.
5.2. End-to-end Performance
In this section, we evaluate the end-to-end performance of Mint, PerColumn, and PerQuery across various workloads. Through these evaluations, we aim to provide comprehensive insights into the practical advantages and trade-offs of Mint compared to PerColumn.

### v2

@@ III-C1 Search and Dynamic Programming Algorithms
$\small\left\{\begin{aligned} cost_{cover}(cover,x_{1})&\qquad\text{when}~i=1,\\ \min_{cvr\subseteq cover}\Big(DP(i-1,cover-cvr)\Big.&\\ \Big.+cost_{cover}(cvr,x_{i})\Big)&\qquad\text{when}~2\leq i\leq|X|.\end{aligned}\right$.
where $cost_{cover}(cvr,x_{i})=\max_{j\in cvr}cost_{idx}(q,x_{i},ek_{i}.j)$ is the cost to cover item subset cvrcvr using index $x_{i}$. The space complexity is $O(2^{k^{\prime}}\cdot|X|)$ for keeping only DP(i1,cover),DP(i,cover),EK(i1,cover)DP(i-1,cover),DP(i,cover),EK(i-1,cover), and $EK(i,cover)$. The time complexity of DP is $O(|X|\cdot 2^{2k^{\prime}})$, bringing the overall time complexity down to $O(k^{\prime}\cdot\sum_{x_{i}\in X}x_{i}.dim+|X|k^{\prime}\log k^{\prime}+|X|\cdot 2^{2k^{\prime}})$. In practice, we set $k^{\prime}$ to a small constant, say $5$, to get acceptable accuracy, and the overall time complexity becomes $O(\sum_{x_{i}\in X}x_{i}.dim+|X|)$, which is linear in the sum of dimensions and number of indexes. Algorithm 1 is the pseudo code for a sample $gt_{sp}(q)$.
In Mint, we use both algorithms in different conditions. When the number of indexes $|X|$ is no more than $3$, we apply the search algorithm. Otherwise, we apply Algorithm 1 for better efficiency.
III-C2 Cost Estimator and Recall Estimator
Here we explain how to estimate cost and recall for the proposed algorithms above. The idea is to use a small sample of training data and historical queries so that we can afford multiple index creations and index scans for cost estimation, as well as full scans for ground truth and recall estimation.
Figure 5: Observed $numDist(q,x_{i},ek_{i})$ using HNSW on four example datasets can be represented by linear models. Similar observation exists for DiskANN [88].
Figure 6: Observed recall using HNSW on four example datasets can be represented by logarithmic models. Similar observation exists for DiskANN [88].
@@ III-C2 Cost Estimator and Recall Estimator
As we will see in Section V, the cost and recall estimators perform reasonably well in practice. The training data size is only 1% of the database and the training time is between 3 and 32 seconds. Note that our algorithms are flexible and can take any cost and recall estimators (Figure 1) if more accurate estimators become available.
IV Configuration Searcher
Configuration Searcher is the main module that searches for the best configuration to solve Problem 1. It explores various configurations, and for each promising configuration, it asks the Query Planner “if these indexes are created, what estimated cost and recall will the best query plan have” (similar to a what-if call in a relational database [17]). Next, we prove its NP-Hardness and propose a beam search algorithm.
IV-A NP-Hardness of Configuration Search
We prove the NP-Hardness of Configuration Search by reducing from the Densest $\mathtt{k}$-Subgraph problem [14, 32]. We replace $\mathtt{k}$ using $\mathtt{g}$ hereafter to distinguish $\mathtt{k}$ from the top-$k$ for queries.
Problem 4 (Densest $\mathtt{g}$-Subgraph).
Given a graph $\mathtt{G}=(\mathtt{V},\mathtt{E})$, find a subgraph $\mathtt{G}_{sub}=(\mathtt{V}_{sub},\mathtt{E}_{sub})$ with exactly $\mathtt{g}$ vertices such that the number of edges in the subgraph is maximal.
@@ IV-A NP-Hardness of Configuration Search
The idea is to map each vertex to an index and each edge to a cost benefit of a query. When two vertices are in $\mathtt{G}_{sub}$, the query can benefit from the corresponding indexes.
@@
(Sketch) Let each vertex $\mathtt{v}^{i}\in\mathtt{V}$ be a single-column index $x^{i}$, each edge $(\mathtt{v}^{i},\mathtt{v}^{j})\in\mathtt{E}$ be a two-column query $q^{i,j}$ in Problem 1, each probability $p^{i,j}$ be the same constant, and each query find the same top-$k$. Then let $\mathtt{g}$ be $\theta_{storage}$, and $\theta_{recall}=100\%$. Finally, let
$latency(q^{i,j},X)=\left\{\begin{aligned} \small k&\quad\text{when}~\mathtt{v}^{i}\in\mathtt{G}_{sub}\textbf{~and~}\mathtt{v}^{j}\in\mathtt{G}_{sub},\\ \small\mathtt{L}&\quad\text{otherwise}.\end{aligned}\right$.
where $\mathtt{L}\gg k$ is a long latency of scanning almost the whole database. This latency function is possible in practice. Specifically, let each $q^{i,j}$ be a vector of all 0 but the $i$-th and $j$-th dimensions are 1. Let ground truth $gt(q^{i,j})=gt^{i}\cup gt^{j}$, where $|gt^{i}|=|gt^{j}|=k/2$. In $gt^{i}$, the vectors’ $i$-th dimension is 1 and other dimensions are 0. Similarly, $gt^{j}$ has vectors with only $j$-th dimension equal to 1. All other dummy vectors are [0.1,,0.1][0.1,...,0.1]. So ground truth vectors have $score=1$ and dummy vectors have $score=0.2$. Plus, $gt^{i}$ rank at the top of $x^{i}$ but $gt^{j}$ rank near the bottom of $x^{i}$ (by scorescore on $\textbf{v}_{i}$), and $gt^{j}$ rank at the top of $x^{j}$ but $gt^{i}$ rank near the bottom of $x^{j}$ (by scorescore on $\textbf{v}_{j}$). As a result, when $\mathtt{v}^{i},\mathtt{v}^{j}\in\mathtt{G}_{sub}$, meaning $x^{i}$ and $x^{j}$ are both picked, the latency is $k$ by retrieving $gt^{i}$ from $x^{i}$ and $gt^{j}$ from $x^{j}$. Otherwise, we must scan other dummy items in latency $\mathtt{L}$.
Now each subgraph $\mathtt{G}_{sub}$ can be mapped to a configuration $X$ and vice versa. The total latency is $(|\mathtt{E}|-|\mathtt{E}_{sub}|)\cdot\mathtt{L}+|\mathtt{E}_{sub}|\cdot k=|\mathtt{E}|\mathtt{L}-|\mathtt{E}_{sub}|(\mathtt{L}-k)$. So maximizing $|\mathtt{E}_{sub}|$ is equivalent to minimizing the total latency and vice versa.
IV-B Beam Search Algorithm
Input: Query workload $W$; Constraints $\theta_{storage}$ and $\theta_{recall}$; Subset difference didi; Seed limit sese; Beam width $b$; Improvement threshold imim; Query Planner.
@@ IV-B Beam Search Algorithm
$Cand(q_{i})\leftarrow\{x~|~x.vid\subseteq q_{i}.vid\text{~and~}|x.vid|\geq|q_{i}.vid|-di\}$
$Candidate\leftarrow\bigcup_{q_{i}}Cand(q_{i})$
$Seed\leftarrow\bigcup_{q_{i}}\{X~|~X\subseteq Cand(q_{i})\text{~and~}|X|\leq se\allowbreak\}$
$Best\leftarrow$ $b$ lowest-cost configurations in SeedSeed that satisfy constraints $\theta_{storage}$ and $\theta_{recall}$
repeat
  $Config\leftarrow\varnothing$
  foreach $X\in Best,x\in Candidate$ do
     $X^{\prime}\leftarrow X\cup\{x\}$
     Get plans for all queries using $X^{\prime}$ from Query Planner and remove unused indexes from $X^{\prime}$
    if $X^{\prime}$ satisfies constraints $\theta_{storage}$ and $\theta_{recall}$ then
        $Config\leftarrow Config\cup X^{\prime}$
     end if
  end foreach
 $Best\leftarrow$ $b$ lowest-cost configurations in ConfigConfig
until Cost improvement $\leq im$
return lowest-cost configuration $X^{*}$ that we have ever seen
@@
We use a beam search algorithm to find optimal configurations as shown in Algorithm 2. We begin by finding a set of candidate indexes from all queries (Line 2 and 2). For each query $q_{i}$, the candidate is a subset whose size is close to $q_{i}.vid$. For instance, when $|q_{i}.vid|=5$ and $di=2$, candidates of $q_{i}$ are subsets with $\geq 3$ columns. Then we create a set of seed configurations from all queries (Line 2). For each query $q_{i}$, the seed configurations are all subsets of $Cand(q_{i})$ with no more than $se$ indexes. Then we keep the best $b$ configurations in Line 2, and perform beam search from Line 2 to Line 2. Within each iteration, we try adding one more index to the existing configuration without violating constraints. The loop stops when the cost improvement $\leq im$.
@@
We implement several optimizations during search. For example, we may search more than $b$ seeds at the beginning to find better configurations if $|Candidate|$ is limited. We also cache the query plans of $(q,X)$ pairs. Instead of calling Query Planner for each $(q,X^{\prime})$ in Line 2, we first identify useful indexes $X^{\prime}_{q}=\{x~|~x\in X^{\prime}\text{~and~}x.vid\subseteq q.vid\text{~and~}|x.vid|\geq|q.vid|-di\}$. Then, if the query plan of $(q,X^{\prime}_{q})$ is cached, we use the cached plan without calling Query Planner. Similarly, to save running time, we cache and pass relevant $ek$ to Query Planner if $ek$ has been computed for a $(q,x)$ pair where $x$ is an index.
V Experiments
We answer the following questions: Q1: What is the quality of indexes and plans generated by Mint for various workloads, compared with state-of-the-art solutions? (Section V-B). Q2: What is the scalability of Mint? How is the runtime of Mint affected by key factors such as query number and storage constraint? (Section V-C).
V-A Setting
We implement the Configuration Searcher and Query Planner in Python, and the underlying indexes in C++. All experiments are conducted on a single machine with an AMD EPYC 7763 64-core processor, 64GB of DDR4 memory, and running Ubuntu 24.04.1 LTS. We repeat each experiment with different random seeds five times and report the mean values. [^2] The code is not open-sourced due to the company policy, but the detailed Algorithms 1 and 2 are sufficient to reproduce all results of the paper. In addition, we release the datasets and workloads in [87].
@@ V-A Setting
AmazonTitle 10M 768
AmazonImage 10M 512
AmazonDescription 10M 768
AmazonFeature 10M 768
AmazonDetail 10M 768
TABLE I: Datasets, each of which is a column (i.e., feature).
Workload Dataset Type # rows # cols # queries Avg. # cols per query
Naïve Semi-Synthetic 1M 3 4 2.00
@@
News Real 0.1M 4 6 2.50
Amazon 10M 5 8 3.25
TABLE II: Workloads.
Datasets and Workloads
To evaluate the performance of Mint, we use a mix of semi-synthetic datasets and real-world datasets (Table I), and five workloads on these datasets (Table II). In each workload, we draw query weights i.i.d. from Uniform(0,1) and normalize them to sum to one.
We use semi-synthetic datasets because most real-world multi-modal datasets have only two columns (text and image), in which case our recommended indexes are the same as the PerColumn baseline. To experiment with more than two columns, we generate semi-synthetic datasets by combining single-feature embedding datasets including GloVe25 [65], GloVe50, GloVe100, GloVe200, SIFT1M [46], Yandex Text-to-Image [73], Deep1M [83], and Music [58]. We sample 1M rows from each single-feature dataset and treat each as a separate column, mimicking the multi-feature scenario. Using these features, we create three workloads: Naïve, BiSimple, and BiComplex. Naïve contains three columns (GloVe100, SIFT1M, and Yandex Text-to-Image), 1M rows, and four manually crafted queries for the case study. BiSimple and BiComplex contain all eight semi-synthetic columns in Table I and 1M rows. Each workload has 12 queries. For each query, each column is independently selected with probability $\mathtt{p}$. Intuitively, a smaller $\mathtt{p}$ generates simpler queries with fewer columns. We use $\mathtt{p}=0.3$ for BiSimple and $\mathtt{p}=0.5$ for BiComplex. For example, a query in BiSimple is on columns GloVe200 and Music, and a query in BiComplex is on GloVe100, SIFT1M, and Deep1M.
One real-world dataset is News, consisting of 0.1M rows and four embeddings extracted from titles, descriptions, contents, and images of BBC News [50]. We get the embeddings for images and titles using the CLIP model [68], which handles images and short text with less than $77$ tokens. For longer text, such as news description and content, we use Nomic-BERT model [61]. The News workload has six randomly-generated queries on the four real embeddings. In each query, column are selected independently with probability $\mathtt{p}=0.5$. One example query is on columns NewsTitle and NewsContent.
We also experiment with a large-scale Amazon [41] dataset including 10M rows and five embeddings to verify the performance gain in Section V-C. Its image embeddings are generated using CLIP, and its title, description, feature, and detail embeddings are generated using Nomic-BERT. It has eight queries. For each query, each column is independently selected with probability $\texttt{p}=0.5$. One example query is on columns AmazonTitle and AmazonFeature.
Figure 7: Using HNSW, (A) Mint outperforms PerColumn on all four workloads, and Mint’s latency is close to the optimal PerQuery’s, although PerQuery uses 33% to 50% more storage. (B) The speedup ranges from $2.1\times$ to $7.2\times$. (C) Mint outperforms PerColumn because Mint has lower cost.
@@
We compare Mint against two baselines: PerColumn and PerQuery. PerColumn means one-index-per-column [77, 86]. It constructs an independent index for each column in the database. Given a workload, each query is answered by the indexes corresponding to the columns involved in the query. PerQuery means one-index-per-query. It creates one index for each query in the workload, so it violates the storage constraints in all workloads we experiment with. However, we still consider it to be a lower bound of latency. All approaches (Mint, PerColumn, and PerQuery) use two widely adopted ANN indexes, HNSW and DiskANN. In Mint, we sample 1% of the data to train the cost estimator and recall estimator (Section III-C2).
@@
We use cosine similarity as the distance measure in all experiments. We measure the latency and the cost as a proxy for latency (Section III-A). Given a workload of queries with probabilities, we measure the probability-weighted average latency and cost across queries.
@@
The storage in bytes is proportional to the number of indexes when a given workload has a fixed number of items, and our maximum degree of HNSW and DiskANN defaults to $16$. So we show the number of indexes as the storage. (Recall in Section II-B that the index graph size is independent of the vectors’ dimensions.)
The default storage constraint is the number of columns because PerColumn requires one index per column. As Table II shows, the storage constraints of Naïve, BiSimple, BiComplex, and News are $3,8,8$, and $4$, respectively. We apply the default storage to PerColumn and Mint for a fair comparison. PerQuery requires one index per query and uses $4,12,12$, and $6$ storage, which violates the constraint but is shown as a reference. In addition, since Mint is designed to be flexible with storage, we evaluate Mint under various storage constraints.
V-B End-to-end Performance
Here we evaluate the end-to-end performance of Mint, PerColumn, and PerQuery across Naïve, BiSimple, BiComplex, and News workloads.